# On the Parameter Identifiability of Partially Observed Linear Causal Models

### v1

@@ A.3 Proof of Theorem 2
Let $\Sigma_{\mathbf{X}}$ be the observed covariance matrix entailed by $A,B,C,D,\Omega_{\mathbf{X}}$ with $\Omega_{\mathbf{L}}=\mathbf{I}$. Let $Q$ be an orthogonal matrix, and
$\hat{A}\coloneqq Q^{T}AQ,\quad\hat{B}\coloneqq Q^{T}B,\quad\hat{C}\coloneqq CQ,\quad\hat{D}\coloneqq D,\quad\hat{\Omega}_{\mathbf{L}}\coloneqq\mathbf{I},\quad\text{and}\quad\hat{\Omega}_{\mathbf{X}}\coloneqq\Omega_{\mathbf{X}}.$
Then, the matrices $\hat{A},\hat{B},\hat{C},\hat{D},\hat{\Omega}_{X}$, and $\hat{\Omega}_{L}$ can also entail the covariance matrix $\Sigma_{\mathbf{X}}$.
@@
Let $M$ and $N$ be matrices defined as in Proposition 1, and similarly for $\hat{M}$ and $\hat{N}$.
$\hat{M}$ $=\left(I-\hat{A}-\hat{B}(I-\hat{D})^{-1}\hat{C}\right)^{-1}$
@@
$\hat{N}$ $=\left((I-\hat{A})\hat{C}^{-1}(I-\hat{D})-\hat{B}\right)^{-1}$
@@
By Proposition 1, the latent covariance matrix $\hat{\Sigma}_{\mathbf{L}}$ is given by
$\hat{\Sigma}_{\mathbf{L}}$ $=\hat{M}^{T}\hat{\Omega}_{\mathbf{L}}\hat{M}+\hat{N}^{T}\hat{\Omega}_{\mathbf{X}}\hat{N}$
$=(Q^{T}M^{T}Q^{-T})(Q^{-1}MQ)+Q^{T}N^{T}\Omega_{\mathbf{X}}NQ$
$=Q^{T}(M^{T}M+N^{T}\Omega_{\mathbf{X}}N)Q$
$=Q^{T}\Sigma_{L}Q.$
By Proposition 1, the observed covariance matrix $\hat{\Sigma}_{\mathbf{X}}$ is given by
$\hat{\Sigma}_{\mathbf{X}}$ $=(I-\hat{D})^{-T}\Big{(}\hat{\Omega}_{\mathbf{X}}+\hat{B}^{T}\hat{\Sigma}_{L}B+\hat{\Omega}_{\mathbf{X}}\hat{N}\hat{B}+\hat{B}^{T}\hat{N}^{T}\hat{\Omega}_{\mathbf{X}}\Big{)}(I-\hat{D})^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}}+(Q^{T}B)^{T}(Q^{T}\Sigma_{L}Q)(Q^{T}B)+\Omega_{\mathbf{X}}(NQ)(Q^{T}B)+(Q^{T}B)^{T}(NQ)^{T}\Omega_{\mathbf{X}}\Big{)}(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}}+B^{T}\Sigma_{L}B+\Omega_{\mathbf{X}}NB+B^{T}N^{T}\Omega_{\mathbf{X}}\Big{)}(I-D)^{-1}$
@@
This indicates that the matrices $\hat{A},\hat{B},\hat{C},\hat{D},\hat{\Omega}_{X}$, and $\hat{\Omega}_{L}$ can also entail the covariance matrix $\Sigma_{\mathbf{X}}$. ∎
Using Lemma 3, we can prove Theorem 2. See 2
@@
Let $U$ be an $(m+n)\times(m+n)$ orthogonal matrix such that: (i) $U_{\mathbf{S}_{1},\mathbf{S}_{1}}=Q$, (ii) the remaining diagonal entries are ones, and (iii) the remaining non-diagonal entries are zeros. Let
$\hat{A}\coloneqq U^{T}AU,\quad\hat{B}\coloneqq U^{T}B,\quad\hat{C}\coloneqq CU,\quad\hat{D}\coloneqq D,\quad\hat{\Omega}_{\mathbf{L}}\coloneqq\mathbf{I},\quad\text{and}\quad\hat{\Omega}_{\mathbf{X}}\coloneqq\Omega_{\mathbf{X}}.$
@@
By construction of $U$, left multiplication of $U^{T}$ on $B$ only affects $B_{\mathbf{S}_{1},*}$; specifically, it is equivalent to replacing $B_{\mathbf{S}_{1},*}$ with $Q^{T}B_{\mathbf{S}_{1},*}$. Furthermore, only the columns of $\mathbf{S}_{5}$ in $B_{\mathbf{S}_{1},*}$ will be affected, because those columns correspond to the measured children of $\mathbf{L}$. Therefore, all entries of $\hat{B}$ are the same as $B$, except that $B_{\mathbf{S}_{1},\mathbf{S}_{5}}$ is replaced with $Q^{T}B_{\mathbf{S}_{1},\mathbf{S}_{5}}$. Similar reasoning shows that all entries of $\hat{C}$ are the same as $C$, except that $C_{\mathbf{S}_{4},\mathbf{S}_{1}}$ is replaced with $C_{\mathbf{S}_{4},\mathbf{S}_{1}}Q$.
Now consider $U^{T}AU$. By the reasoning above, left multiplication of $U^{T}$ on $A$ only is equivalent to replacing $A_{\mathbf{S}_{1},\mathbf{S}_{3}}$ with $Q^{T}A_{\mathbf{S}_{1},\mathbf{S}_{3}}$. Further right multiplication of $U$ on $U^{T}A$ is equivalent to replacing $(U^{T}A)_{\mathbf{S}_{2},\mathbf{S}_{1}}$ with $(U^{T}A)_{\mathbf{S}_{2},\mathbf{S}_{1}}Q$. Since $\mathbf{S}_{1}$, $\mathbf{S}_{2}$, and $\mathbf{S}_{3}$ are mutually disjoint, all entries of $\hat{A}=U^{T}AU$ are the same as $A$, except that $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}$ and $A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$ are replaced with $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}Q$ and $Q^{T}A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$, respectively.
Hence, for matrices $A,B,C$, and $D$, suppose we replace $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}$, $A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$, $C_{\mathbf{S}_{4},\mathbf{S}_{1}}$, and $B_{\mathbf{S}_{1},\mathbf{S}_{5}}$ with $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}Q$, $Q^{T}A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$, $C_{\mathbf{S}_{4},\mathbf{S}_{1}}Q$, and $Q^{T}B_{\mathbf{S}_{1},\mathbf{S}_{5}}$, respectively. By the reasoning above, this is equivalent to replacing $A$, $B$, $C$, and $D$ with $\hat{A}$, $\hat{B}$, $\hat{C}$, and $\hat{D}$, respectively, which can entail the same covariance matrix $\Sigma_{\mathbf{X}}$. ∎
A.4 Proof of Corollary 1 and Corollary 2
@@ A.5 Proof of Proposition 1
Since matrices $A$ and $D$ are invertible, using the formula of $2\times 2$ block matrix inversion [21, Chapter 0.7], we obtain
@@
$(I-F)^{-T}\Omega=\begin{pmatrix}M^{T}\Omega_{\mathbf{L}}&-M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}\\ -(I-D)^{-T}B^{T}M^{T}\Omega_{\mathbf{L}}&(I-D)^{-T}\Omega_{\mathbf{X}}+(I-D)^{-T}B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}\end{pmatrix}.$
@@
$\Sigma_{L}$ $=M^{T}\Omega_{\mathbf{L}}M+M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}(I-D)^{-1}CM$
$=M^{T}\Omega_{\mathbf{L}}M+N^{T}\Omega_{\mathbf{X}}N$
@@
$\Sigma_{X}$ $=(I-D)^{-T}B^{T}M^{T}\Omega_{\mathbf{L}}MB(I-D)^{-1}+(I-D)^{-T}\Omega_{\mathbf{X}}(I-D)^{-1}$
$\qquad+(I-D)^{-1}\Omega_{\mathbf{X}}(I-D)^{-1}CMB(I-D)^{-1}+(I-D)^{-T}B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}(I-D)^{-1}$
$\qquad+(I-D)^{-T}B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}(I-D)^{-1}CMB(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}}+B^{T}M^{T}\Omega_{\mathbf{L}}MB+\Omega_{\mathbf{X}}(I-D)^{-1}CMB+B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}$
$\qquad\qquad\qquad\qquad+B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\mathbf{X}}(I-D)^{-1}CMB\Big{)}(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\mathbf{X}}+B^{T}\Sigma_{L}B+\Omega_{\mathbf{X}}NB+B^{T}N^{T}\Omega_{\mathbf{X}}\Big{)}(I-D)^{-1}.$
A.6 Computational Cost of Checking Whether the Conditions in Theorem 1 Hold
@@
A.7 In Practice, What If the Conditions Do not Hold?
@@ B.2 Definition of Pure Children
Definition 7 (Pure Children [17]).
$\mathbf{Y}$ are pure children of variables $\mathbf{X}$ in graph $\mathcal{G}$, iff $\text{Pa}_{\mathcal{G}}(\mathbf{Y})=\cup_{\mathsf{Y_{i}}\in\mathbf{Y}}\text{Pa}_{\mathcal{G}}(\mathsf{Y_{i}})=\mathbf{X}$ and $\mathbf{X}\cap\mathbf{Y}=\emptyset$. We denote the pure children of $\mathbf{X}$ in $\mathcal{G}$ by $\text{PCh}_{\mathcal{G}}(\mathbf{X})$.
B.3 Definition of Rank-invariant Graph Operator

### v2

@@ A.5 Proof of Theorem 4
Let $\Sigma_{\mathbf{X}}$ be the observed covariance matrix entailed by $F_{\mathbf{L}\mathbf{L}},F_{\mathbf{L}\mathbf{X}},F_{\mathbf{X}\mathbf{L}},F_{\mathbf{X}\mathbf{X}},\Omega_{\epsilon_{\mathbf{L}}},\Omega_{\epsilon_{\mathbf{X}}}$. Let $Q$ be an orthogonal matrix, and
$\tilde{F}_{\mathbf{L}\mathbf{L}}=Q^{T}F_{\mathbf{L}\mathbf{L}}Q,~{}\tilde{F}_{\mathbf{L}\mathbf{X}}=Q^{T}F_{\mathbf{L}\mathbf{X}},~{}\tilde{F}_{\mathbf{X}\mathbf{L}}=F_{\mathbf{X}\mathbf{L}}Q,~{}\tilde{F}_{\mathbf{X}\mathbf{X}}=F_{\mathbf{X}\mathbf{X}},~{}\tilde{\Omega}_{\epsilon_{\mathbf{L}}}=Q^{T}\Omega_{\epsilon_{\mathbf{L}}}Q,~{}\text{and}~{}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}=\Omega_{\epsilon_{\mathbf{X}}}.$
Then, the matrices $\tilde{F}_{\mathbf{L}\mathbf{L}},\tilde{F}_{\mathbf{L}\mathbf{X}},\tilde{F}_{\mathbf{X}\mathbf{L}},\tilde{F}_{\mathbf{X}\mathbf{X}},\tilde{\Omega}_{\epsilon_{\mathbf{L}}},\tilde{\Omega}_{\epsilon_{\mathbf{X}}}$ can also entail the covariance matrix $\Sigma_{\mathbf{X}}$.
@@
[…]
$\tilde{M}$ $=\left(I-\tilde{A}-\tilde{B}(I-\tilde{D})^{-1}\tilde{C}\right)^{-1}$
@@
$\tilde{N}$ $=\left((I-\tilde{A})\tilde{C}^{-1}(I-\tilde{D})-\tilde{B}\right)^{-1}$
@@
By Proposition 1, the latent covariance matrix $\tilde{\Sigma}_{\mathbf{L}}$ is given by
$\tilde{\Sigma}_{\mathbf{L}}$ $=\tilde{M}^{T}\tilde{\Omega}_{\epsilon_{\mathbf{L}}}\tilde{M}+\tilde{N}^{T}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\tilde{N}$
$=(Q^{T}M^{T}Q^{-T})(Q^{T}\Omega_{\epsilon_{\mathbf{L}}}Q)(Q^{-1}MQ)+Q^{T}N^{T}\Omega_{\epsilon_{\mathbf{X}}}NQ$
$=Q^{T}(M^{T}\Omega_{\epsilon_{\mathbf{L}}}M+N^{T}\Omega_{\epsilon_{\mathbf{X}}}N)Q$
$=Q^{T}\Sigma_{\mathbf{L}}Q.$
By Proposition 1, the observed covariance matrix $\tilde{\Sigma}_{\mathbf{X}}$ is given by
$\tilde{\Sigma}_{\mathbf{X}}$ $=(I-\tilde{D})^{-T}\Big{(}\tilde{B}^{T}\tilde{\Sigma}_{\mathbf{L}}B+\tilde{\Omega}_{\epsilon_{\mathbf{X}}}+\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\tilde{N}\tilde{B}+\tilde{B}^{T}\tilde{N}^{T}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\Big{)}(I-\tilde{D})^{-1}$
$=(I-D)^{-T}\Big{(}(Q^{T}B)^{T}(Q^{T}\Sigma_{\mathbf{L}}Q)(Q^{T}B)+\Omega_{\epsilon_{\mathbf{X}}}+\Omega_{\epsilon_{\mathbf{X}}}(NQ)(Q^{T}B)+(Q^{T}B)^{T}(NQ)^{T}\Omega_{\epsilon_{\mathbf{X}}}\Big{)}(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}B^{T}\Sigma_{\mathbf{L}}B+\Omega_{\epsilon_{\mathbf{X}}}+\Omega_{\epsilon_{\mathbf{X}}}NB+B^{T}N^{T}\Omega_{\epsilon_{\mathbf{X}}}\Big{)}(I-D)^{-1}$
@@
This indicates that the matrices $\tilde{A},\tilde{B},\tilde{C},\tilde{D},\tilde{\Omega}_{\epsilon_{\mathbf{L}}}$ and $\tilde{\Omega}_{\epsilon_{\mathbf{X}}}$ can also entail the covariance matrix $\Sigma_{\mathbf{X}}$. ∎
Using Lemma 3, we can prove Theorem 4. See 4
@@
[…]
Let $U$ be an $m\times m$ orthogonal matrix such that: (i) $U_{\mathbf{S}_{1},\mathbf{S}_{1}}=Q$, (ii) the remaining diagonal entries are ones, and (iii) the remaining non-diagonal entries are zeros. Let
$\tilde{A}\coloneqq U^{T}AU,\quad\tilde{B}\coloneqq U^{T}B,\quad\tilde{C}\coloneqq CU,\quad\tilde{D}\coloneqq D,\quad\tilde{\Omega}_{\epsilon_{\mathbf{L}}}\coloneqq U^{T}\Omega_{\epsilon_{\mathbf{L}}}U=I,\quad\text{and}\quad\tilde{\Omega}_{\epsilon_{\mathbf{X}}}\coloneqq\Omega_{\epsilon_{\mathbf{X}}}.$
@@
By construction of $U$, left multiplication of $U^{T}$ on $B$ only affects $B_{\mathbf{S}_{1},*}$; specifically, it is equivalent to replacing $B_{\mathbf{S}_{1},*}$ with $Q^{T}B_{\mathbf{S}_{1},*}$. Furthermore, only the columns of $\mathbf{S}_{5}$ in $B_{\mathbf{S}_{1},*}$ will be affected, because those columns correspond to the measured children of $\mathbf{L}$. Therefore, all entries of $\tilde{B}$ are the same as $B$, except that $B_{\mathbf{S}_{1},\mathbf{S}_{5}}$ is replaced with $Q^{T}B_{\mathbf{S}_{1},\mathbf{S}_{5}}$. Similar reasoning shows that all entries of $\tilde{C}$ are the same as $C$, except that $C_{\mathbf{S}_{4},\mathbf{S}_{1}}$ is replaced with $C_{\mathbf{S}_{4},\mathbf{S}_{1}}Q$.
Now consider $U^{T}AU$. By the reasoning above, left multiplication of $U^{T}$ on $A$ only is equivalent to replacing $A_{\mathbf{S}_{1},\mathbf{S}_{3}}$ with $Q^{T}A_{\mathbf{S}_{1},\mathbf{S}_{3}}$. Further right multiplication of $U$ on $U^{T}A$ is equivalent to replacing $(U^{T}A)_{\mathbf{S}_{2},\mathbf{S}_{1}}$ with $(U^{T}A)_{\mathbf{S}_{2},\mathbf{S}_{1}}Q$. Since $\mathbf{S}_{1}$, $\mathbf{S}_{2}$, and $\mathbf{S}_{3}$ are mutually disjoint, all entries of $\tilde{A}=U^{T}AU$ are the same as $A$, except that $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}$ and $A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$ are replaced with $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}Q$ and $Q^{T}A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$, respectively.
Hence, for matrices $A,B,C$, and $D$, suppose we replace $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}$, $A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$, $C_{\mathbf{S}_{4},\mathbf{S}_{1}}$, and $B_{\mathbf{S}_{1},\mathbf{S}_{5}}$ with $A_{\mathbf{S}_{2},{\mathbf{S}}_{1}}Q$, $Q^{T}A_{\mathbf{S}_{1},{\mathbf{S}}_{3}}$, $C_{\mathbf{S}_{4},\mathbf{S}_{1}}Q$, and $Q^{T}B_{\mathbf{S}_{1},\mathbf{S}_{5}}$, respectively. By the reasoning above, this is equivalent to replacing $A$, $B$, $C$, and $D$ with $\tilde{A}$, $\tilde{B}$, $\tilde{C}$, and $\tilde{D}$, respectively, which (generically) share the same support and entail the same covariance matrix $\Sigma_{\mathbf{X}}$. ∎
A.6 Proof of Corollary 1 and Corollary 2
@@ A.7 Proof of Proposition 1
[…]
Since matrices $A$ and $D$ are invertible, using the formula of $2\times 2$ block matrix inversion [22, Chapter 0.7], we obtain
@@
$(I-F)^{-T}\Omega=\begin{pmatrix}M^{T}\Omega_{\epsilon_{\mathbf{L}}}&-M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}\\ -(I-D)^{-T}B^{T}M^{T}\Omega_{\epsilon_{\mathbf{L}}}&(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}+(I-D)^{-T}B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}\end{pmatrix}.$
@@
$\Sigma_{L}$ $=M^{T}\Omega_{\epsilon_{\mathbf{L}}}M+M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}CM$
$=M^{T}\Omega_{\epsilon_{\mathbf{L}}}M+N^{T}\Omega_{\epsilon_{\mathbf{X}}}N$
@@
$\Sigma_{X}$ $=(I-D)^{-T}B^{T}M^{T}\Omega_{\epsilon_{\mathbf{L}}}MB(I-D)^{-1}+(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}$
$\qquad+(I-D)^{-1}\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}CMB(I-D)^{-1}+(I-D)^{-T}B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}$
$\qquad+(I-D)^{-T}B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}CMB(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\epsilon_{\mathbf{X}}}+B^{T}M^{T}\Omega_{\epsilon_{\mathbf{L}}}MB+\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}CMB+B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}$
$\qquad\qquad\qquad\qquad+B^{T}M^{T}C^{T}(I-D)^{-T}\Omega_{\epsilon_{\mathbf{X}}}(I-D)^{-1}CMB\Big{)}(I-D)^{-1}$
$=(I-D)^{-T}\Big{(}\Omega_{\epsilon_{\mathbf{X}}}+B^{T}\Sigma_{\mathbf{L}}B+\Omega_{\epsilon_{\mathbf{X}}}NB+B^{T}N^{T}\Omega_{\epsilon_{\mathbf{X}}}\Big{)}(I-D)^{-1}.$
We now discuss how $M$ and $N$ defined in Proposition 1 are invertible. Note that matrices $I-D$ and $I-F$ are invertible because structure $\mathcal{G}$ is acyclic. This implies $\det(I-F)\neq 0$ and $\det(I-D)\neq 0$. Define
$U=\begin{pmatrix}I&0\\ \\ -(I-D)^{-1}C&I\end{pmatrix},$
which implies
$(I-F)U=\begin{pmatrix}M&B\\ \\ 0&I-D\end{pmatrix}$
and thus
$\det((I-F)U)=\det(M)\det(I-D).$
Since $\det(U)=1$ and $\det(I-F)\neq 0$, we have
$\det((I-F)U)=\det(I-F)\det(U)\neq 0.$
By the statement above and $\det(I-D)\neq 0$, we have
$\det(M)=\frac{\det((I-F)U)}{\det(I-D)}\neq 0,$
which implies that $M$ is invertible. Similar reasoning can be used to show that $N$ is invertible.
A.8 Computational Cost of Checking Whether the Conditions in Theorem 3 Hold
@@
A.9 In Practice, What If the Conditions Do not Hold?
@@ B.2 Definition of Pure Children
Definition 6 (Parents, Children, and Descendants of a Set of Nodes [18]).
For a set of nodes $\mathbf{X}$ in $\mathcal{G}$, we have $\text{Ch}_{\mathcal{G}}(\mathbf{X})=\cup_{\mathsf{X}\in\mathbf{X}}\text{Ch}_{\mathcal{G}}(\mathsf{X})$, $\text{Pa}_{\mathcal{G}}(\mathbf{X})=\cup_{\mathsf{X}\in\mathbf{X}}\text{Pa}_{\mathcal{G}}(\mathsf{X})$, and $\text{De}_{\mathcal{G}}(\mathbf{X})=\cup_{\mathsf{X}\in\mathbf{X}}\text{De}_{\mathcal{G}}(\mathsf{X})$.
Definition 7 (Pure Children of a Set of Nodes [18]).
$\mathbf{Y}$ are pure children of a set of nodes $\mathbf{X}$ in graph $\mathcal{G}$, i.e., $\mathbf{Y}\in\text{PCh}_{\mathcal{G}}(\mathbf{X})$, iff all of the following hold: (i) $\mathbf{X}\cap\mathbf{Y}=\emptyset$, (ii) $\mathbf{Y}\subseteq\text{Ch}_{\mathcal{G}}(\mathbf{X})$, (iii) $\text{Pa}_{\mathcal{G}}(\mathbf{Y})=\mathbf{X}$, and (iv) $\text{De}_{\mathcal{G}}(\mathbf{Y})\cap\mathbf{X}=\emptyset$.
B.3 Definition of Neighbor and MEC
Definition 8 (Neighbor).
In $\mathcal{G}$, nodes $\mathsf{X}$ and $\mathsf{Y}$ are neighbor of each other iff there exist an edge from $\mathsf{X}$ to $\mathsf{Y}$ or an edge from $\mathsf{Y}$ to $\mathsf{X}$.
Definition 9 (Markov Equivalence Class (MEC)).
Two DAGS $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ belong to the same MEC, iif they share the same skeleton and v-structures.
B.4 Definition of Rank-invariant Graph Operator